OpenStudy (jellybot23):

A roller coaster weighing 8,000N descends a 60 degrees slope. What is the roller coaster's rate of acceleration (ignoring friction and air resistance). I know the answer to this and the steps to the answer, but I don't understand how exactly to do it. I wanna go into a physics career, so I need the steps really broken down and explained

OpenStudy (michele_laino):

we have to find the force along the direction of the slope, which is acting on the roller coaster, like below: 

OpenStudy (michele_laino):

from the drawing above, the force acting on the roller coaster, has the subsequent magnitude: \(mg \sin \theta\)

OpenStudy (michele_laino):

so, the corresponding acceleration is: \[a = \frac{F}{m} = \frac{{mg\sin \theta }}{m} = g\sin \theta \] wher \(g=9.81\) m/sec^2

OpenStudy (michele_laino):

where*

OpenStudy (jellybot23):

Oh okay, wow that clears it up a lot! So it is basically just substituting in for Newton's 2nd law? correct?

OpenStudy (michele_laino):

yes! we have to apply the famous equation: \(F=m \cdot a\)

OpenStudy (jellybot23):

Okay awesome! Thank you so much! That was a lot simpler than I was making it haha :)

OpenStudy (michele_laino):

:)